The amount of material used in making a custom sail for a sailboat is normally distributed with a standard deviation of 64 square feet. For a random sample of 15 sails, the mean amount of material used is 912 square feet. Find a 90% confidence interval for the population mean amount of material used in a custom sail.

  1. Point estimate value = 
  2. Standard error (approximate to four decimal points) = 
  3. Margin of error (approximate to four decimal points) = 
  4. Tabulated value (approximate to four decimal points) = 
  5. Upper limit (approximate to four decimal points) = 
  6. Lower limit (approximate to four decimal points) = 
1
Expert's answer
2022-05-16T17:54:32-0400

1. Point estimate value = 


"\bar{x}=912/15=60.8"



2. Standard error (approximate to four decimal points) = 


"SE=\sigma/\sqrt{n}=64/\sqrt{15}=16.5247"

3. Margin of error (approximate to four decimal points) = 


"M=z_{\alpha/2}\times SE=1.6449\times\dfrac{64}{\sqrt{15}}=27.1815"

4. Tabulated value (approximate to four decimal points) = 


"z_{\alpha/2}=1.6449"


5. Upper limit (approximate to four decimal points) = 


"60.8+16.5247=77.3247"

6. Lower limit (approximate to four decimal points) = 


"60.8-16.5247=44.2753"